# Probing the nature of gravity in the low-acceleration limit: wide binaries of extreme separations with perspective effects

### v1

@@ 6 Discussion and Conclusion
Our study indicates that the probe of gravity in the low-acceleration limit with $s>30$ kau will be quite limited even in the future, when based only on conventional statistical methods as considered here. This is because the number of isolated wide binaries in the solar neighborhood may not be sufficient to obtain a desired precision of $\gamma_{g}$. This difficulty may be overcome by the Bayesian 3D methodology (K.-H. Chae, 2025) with accurate measured radial velocities in the future, because it can return a precise determination of $\gamma_{g}$ even with dozens of wide binaries once high-precision radial velocities become available.
This work was supported by the National Research Foundation of Korea (NRF-2022R1A2C1092306).

### v2

@@ 6 Discussion and Conclusion
Our study indicates that the probe of gravity in the low-acceleration limit with $s>30$ kau will be quite limited even in the future, when based only on conventional statistical methods as considered here. This is because the number of isolated wide binaries in the solar neighborhood may not be sufficient to obtain a desired precision of $\gamma_{g}$. This difficulty may be overcome by the Bayesian 3D methodology (K.-H. Chae, 2025a, b) with accurate measured radial velocities in the future, because it can return a precise determination of $\gamma_{g}$ even with dozens of wide binaries once high-precision radial velocities become available.
[…]
We thank the reviewer for the constructive feedback and valuable suggestions. This work was supported by the National Research Foundation of Korea (NRF-2022R1A2C1092306). YT also acknowledges Taiwan National Science and Technology Council (NSTC) grant 114-2112-M-008-024-MY3.
@@ Appendix B Results with varied ruwe limits
Appendix C Robustness Test against Contamination from Photometric Binaries
In this appendix, we test the robustness of our main results against potential contamination from unresolved close binaries, often referred to as photometric binaries. As noted in studies such as V. Belokurov et al. (2020); X. Hernandez (2023), these systems can be identified on a Color-Magnitude Diagram (CMD) as they typically form a sequence located above the single-star main sequence due to photometric blending. While our primary sample selection, which includes a ${\tt ruwe}<1.2$ cut, is effective at removing many unresolved multiples, some may remain and potentially bias the kinematic analysis.
To address this, we implement a more conservative CMD selection strategy inspired by the methodology of X. Hernandez (2023). We define a main-sequence ridge line by connecting the points (BP-RP, $M_{G}$) = (0.7, 4.7) and (2.2, 9.7). We then construct a selection band around this line, designed to follow the observed main sequence more closely. Specifically, we exclude any binary system where either component lies more than 0.8 magnitudes above or below this ridge line. While the lower boundary is kept at 0.8 magnitudes below, the upper boundary is further adjusted as a function of color to account for the main sequence’s intrinsic curvature, ensuring a precise exclusion of the photometric binary population as illustrated in Figure 13. We note that there can be single stars above the upper boundary and the cut is not designed to selectively remove only photometric binaries. The application of this stricter cut reduces our sample size from 8,833 to 7,800 pairs.
The results of the acceleration-plane analysis using this cleaner sample are presented in Figure 14. When compared with the analysis of our nominal sample (Figure 6), the magnitude of the observed gravitational anomaly is slightly attenuated, but the results remain quantitatively consistent and the statistical significance of the deviation from the Newtonian prediction is largely unchanged. We therefore conclude that our primary findings are robust, and the observed anomaly cannot be attributed to kinematic contamination from the population of photometric binaries removed in this test.
Figure 13: We apply additional cuts in the CMD to exclude the photometric binary sequence lying above the main sequence (Figure 1), which is produced by unresolved close binaries. After this cleaning, 7,800 binaries remain from the initial 8,833 pairs.
Figure 14: Acceleration-plane analysis of our sample obtained by removing the potential contamination from unresolved close binaries as in Figure 13. Compared with Figure 6, there is no significant difference.
Appendix D Results for binaries with both radial velocities available
As mentioned before, in our analysis so far, we considered binaries with both radial velocities available, or one of the radial velocities available. In this appendix, we present the acceleration-plane analysis for binaries with both radial velocities available. In Figure 15, we see that there is no tangible difference from Figure 6.
Figure 15: Acceleration-plane analysis for binaries with both radial velocities available. The number of binaries is 7,250, slightly smaller than our nominal sample 8,833. There is no appreciable difference from Figure 6.